Show that the MacLaurin's series for f (z) =sin z is given by

z z 3 s "' (- I)" z211+J

f(z) =z--+- ....+L ( ) +...


1
Expert's answer
2021-09-27T15:01:59-0400

"\displaystyle
\text{The maclaurin's series is given by}\\
f(z) = f(0) + zf'(0) + \frac{z^2}{2}f''(0) + \frac{z^3}{6}f'''(0)+\frac{z^4}{24}f^{iv}(0)+\frac{z^5}{120}f^{v}(0)+\cdots\\
\text{Where}\\
f(0) = sin0=0\\
f'(z) = cosz, \qquad f'(0) = 1\\
f''(z) = -sinz, \qquad f''(0) =0\\
f'''(z) = -cosz, \qquad f'''(0) =-1\\
f^{iv}(z) = sinz, \qquad f^{iv}(0) =0\\
f^{v}(z) = cosz, \qquad f^{v}(0) =1\\
\text{Hence, the maclaurin's series is given by}\\
f(z) = z-\frac{z^3}{6}+\frac{z^5}{120}+ \cdots"